OpenStudy (anonymous):

Please Help....... Calculate the percent composition of (NH4)2CO3. The correct percentages are 1. 29.15% N, 8.407% H, 12.50% C, 49.94% O. 2. 32.74% N, 6.850% H, 15.67% C, 44.74% O. 3. 25.65% N, 5.670% H, 12.50% C, 46.18% O. 4. 39.15% N, 8.407% H, 12.50% C, 39.94% O.

OpenStudy (cyber405):

Percent Composition=Mass due to specific component/Total molar mass of compound Then multiply it to 100... For example i calculate the percent composition of C on CO2: Mass due to C = 12.01 g/mol Molar mass of compound= 12.01+2*(16)=44.01 g/mol Percent Composition of C = (12.01/44.01)*100 = 27.29 %

OpenStudy (anonymous):

@ cyber405 is right

OpenStudy (toxicsugar22):

help me please

OpenStudy (anonymous):

in which one

OpenStudy (anonymous):

12/96 then times 100 that is 12.5 so 2 IS WRONG 48/96 GIVES ME 50% FOR OXYGEN OR 49.94% AS IN THE ANSWER SO 1 IS CORRECT you know why i took oxygen because it has different values in all of the ANSWER CHOICES